OpenStudy (anonymous):

Two studies were completed in California. One study in northern California involved 1,000 patients; 74% of them experienced flulike symptoms during the month of December. The other study, in southern California, involved 500 patients; 34% of them experienced flulike symptoms during the same month. Which study has the smallest margin of error for a 98% confidence interval?

OpenStudy (anonymous):

The northern California study with a margin of error of 3.2%. The southern California study with a margin of error of 3.2%. The northern California study with a margin of error of 4.9%. The southern California study with a margin of error of 4.9%.

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

Loser I dont get this that much

OpenStudy (anonymous):

@vera_ewing

OpenStudy (anonymous):

I think its southern california

OpenStudy (anonymous):

Just write something and I will give you a medal @Loser66

OpenStudy (anonymous):

What ever you want

OpenStudy (loser66):

hehhehe... I don't want medal, just interesting on the field I never know before. Here is the answer: the first one.

OpenStudy (loser66):

I made a short research about it. I learn how to find the answer out, not guess!

OpenStudy (anonymous):

Help me though

OpenStudy (loser66):

if the confidence is 98% , then z =2.33 for the NOrth one the sample is 74%*1000= 740 hence p = 740/1000=0.74

OpenStudy (loser66):

Apply the formula margin error = \(z*\sqrt{\dfrac{p(1-p)}{population}}=2.33*\sqrt\dfrac{0.74(1-0.74}{1000}\ )

OpenStudy (loser66):

=3.23% the net is so bad.

OpenStudy (anonymous):

Thanks